# LoRI: Reducing Cross-Task Interference in Multi-Task Low-Rank Adaptation

### v1

@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023) have transformed deep learning, showcasing remarkable capabilities across various domains. However, their deployment remains computationally demanding, particularly when fine-tuning is required to adapt to downstream tasks or align with human preferences. To mitigate the high resource costs, researchers have developed a range of parameter-efficient fine-tuning (PEFT) techniques. Among these techniques, LoRA has gained widespread adoption. Nevertheless, LoRA still introduces notable memory overhead, particularly in large-scale models. Consequently, recent research has focused on further optimizing LoRA by reducing the number of trainable parameters without compromising performance (Kopiczko et al., 2023; Ding et al., 2023; Zhang et al., 2023b).
@@
Recent studies (Yu et al., 2024; Panda et al., 2024) have shown that delta parameters – the differences between fine-tuned and pretrained model weights – exhibit significant redundancy. Motivated by the effectiveness of random projections (Aghajanyan et al., 2020; Lu et al., 2022; Zhang et al., 2023b; Tian et al., 2024) and the observed redundancy in delta parameters, we propose LoRA with Reduced Interference (LoRI). LoRI keeps the low-rank matrices $A$ fixed as random projections, while training the matrices $B$ using task-specific sparse masks. To retain the most critical elements of $B$, LoRI performs a calibration process to extract sparse masks by selecting the highest-magnitude elements across all layers and projections. As shown in Figure 1(a), LoRI maintains performance even with 90% sparsity in $B$ while keeping $A$ frozen. This demonstrates that adaptation does not require updating $A$, and that $B$ has considerable redundancy. By applying more constrained updates than LoRA, LoRI significantly reduces the number of trainable parameters while better preserving the pretrained model’s knowledge during adaptation.
@@
Figure 1: (a) Varying sparsity ratios in matrices $B$ while freezing $A$. Performance remains stable even at 90% sparsity in matrices $B$. (b) Merging three adapters via weighted averaging. LoRA suffers degradation due to parameter interference, while LoRI preserves task performance. (c) Continual learning from Safety to NLU. LoRA suffers from catastrophic forgetting, while LoRI retains safety alignment. Results for NLU are averaged over eight tasks. GSM8K accuracy (Math) and HumanEval pass@10 (Code) are reported individually. Base model: Llama-3-8B, rank $r=32$.
@@
Multi-task learning is essential for enabling versatile models with multi-task capabilities, which is traditionally performed via joint training on a combination of task-specific datasets (Caruana, 1997; Sener & Koltun, 2018). However, training large models on this data mixture is prohibitively expensive in terms of time and compute. Model merging is a training-free alternative for building powerful models by combining existing ones (Ilharco et al., 2022; Yadav et al., 2023; Yu et al., 2024). This approach is well-suited for merging LoRA adapters to enable multi-task capabilities within a single LoRA (Wang et al., 2024a; Prabhakar et al., 2024). However, as shown in Figure 1(b), directly merging heterogeneous LoRAs often results in parameter interference, leading to degraded performance in the merged LoRA compared to single-task LoRAs. Additionally, many existing merging methods require trial-and-error to identify the optimal method for a specific combination of tasks. LoRI tackles these challenges by enabling adapter merging without manual selection of merging methods. By using fixed, randomly initialized projection $A$, LoRI maps task-specific adapters into approximately orthogonal subspaces, thereby reducing interference when merging multiple LoRIs.
@@
Beyond multi-tasking, safety-critical scenarios require that each newly introduced adapter enhances model capabilities while preserving the safety alignment of the pretrained base model (Qi et al., 2023). LoRI provides a lightweight continual learning approach for adapting models while preserving safety, where training is performed sequentially across tasks (Lopez-Paz & Ranzato, 2017; Wu et al., 2022; Ouyang et al., 2022). The strategy involves first fine-tuning an adapter on safety data to establish alignment, followed by separate adaptation to each downstream task. However, as illustrated in Figure 1(c), continual learning often leads to catastrophic forgetting (Li & Hoiem, 2017; Dong et al., 2023; Luo et al., 2023), wherein the adaptation to new tasks substantially compromises previously acquired knowledge. LoRI mitigates forgetting by leveraging the sparsity of matrices $B$ through task-specific masks. This isolation of parameter updates across tasks facilitates continual learning with minimal interference, preserving both safety and task effectiveness.
@@ 2.1 Freezing Low-Rank Projections with Sparse Masking
LoRA (Hu et al., 2021) fine-tunes a weight update matrix as a product of two low-rank matrices to adapt LLMs to new tasks. Formally, for a specific task $t$, given a pretrained weight matrix $W\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, the weight update $\Delta_{t}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ is constrained to a low-rank decomposition:
$h=xW+x\Delta_{t}=xW+xA_{t}B_{t}.$ (1)
@@
Typically, the low-rank projection matrix $A_{t}$ and the low-rank expansion matrix $B_{t}$ are updated via gradient descent. Matrix $A_{t}$ is usually initialized with a random Gaussian distribution, while matrix $B_{t}$ is initialized to zero, ensuring that $\Delta_{t}=0$ at the start of training. However, in LoRI, we fix $A_{t}$ as a random projection, meaning that the model only learns how to combine the fixed subspace via $B_{t}$. By freezing $A_{t}$, we eliminate the need to store its gradients and optimizer states, thereby reducing memory consumption. During inference, similar to LoRA, LoRI merges the low-rank updates by adding $A_{t}B_{t}$ to $W$, ensuring no additional inference latency compared to full fine-tuning.
@@
Figure 2: Overview of the proposed LoRI method. (a) LoRI freezes the projection matrices $A_{t}$ and sparsely updates $B_{t}$ using task-specific masks. (b) LoRI enables adapter merging of multiple task-specific adapters with reduced parameter interference. (c) LoRI builds safety adapters by continual learning with reduced catastrophic forgetting.
@@
Here, $\mathbb{I}(\cdot)$ denotes the indicator function applied element-wise. This ensures that only the top-$(1-s)\%$ of parameters (by magnitude) across all layers and projections are retained. It is well established that the importance of projection matrices varies significantly across different layers and modules (Zhang et al., 2023a; d; Kopiczko et al., 2023). Our masking strategy enables global comparison of parameters and facilitates effective allocation of the parameter budget determined by the sparsity ratio. Notably, the masks for each task $t$ are calibrated only once and can be reused as needed.
@@ 2.2 Reducing Interference in Adapter Merging via Orthogonality
A central challenge in adapter merging is parameter interference, where combining multiple adapters leads to degraded performance due to conflicting updates. LoRI mitigates this issue by leveraging orthogonality in the representation space, achieved by freezing the projection matrices $A_{t}$ as random matrices. This leads to the following property:
@@
Let $A_{s},A_{t}\in\mathbb{R}^{d_{\text{in}}\times r}$ be independent random matrices with entries drawn from a zero-mean, unit-variance Gaussian distribution, where $r\ll d_{\text{in}}$. Let $\Delta_{s}=A_{s}(B_{s}\odot M_{s})$ and $\Delta_{t}=A_{t}(B_{t}\odot M_{t})$ be two LoRI adapters. Then, in high-dimensional settings, the adapters are approximately orthogonal with high probability: $\langle\Delta_{s},\Delta_{t}\rangle\approx 0$.
The proof of this property is provided in Appendix C and Lemma 5 of Goldstein & Studer (2018). This property implies that the LoRIs $\Delta_{t}=A_{t}(B_{t}\odot M_{t})$ for different tasks reside in approximately orthogonal subspaces. Therefore, the expected interference between LoRIs is negligible, preserving task-specific performance in the merged model.
Merging LoRI Adapters.
@@
Given a set of trained LoRI adapters $\{\Delta_{1},\Delta_{2},\dots,\Delta_{T}\}$, the goal is to construct a unified adapter $\Delta_{\text{merge}}$ that combines knowledge from all tasks with minimal interference, as illustrated in Figure 2(b). We describe two merging methods: concatenated merging (weighted averaging) and linear merging (Task Arithmetic) (Ilharco et al., 2022), both of which exploit the approximate orthogonality of LoRIs.
Concatenated Merging (Weighted Averaging). Define the concatenated matrices as:
$A^{\prime}=[\alpha_{1}A_{1}\ \alpha_{2}A_{2}\ \dots\ \alpha_{T}A_{T}],\quad B^{\prime}=\left[(B_{1}\odot M_{1})^{\top},(B_{2}\odot M_{2})^{\top},\dots,(B_{T}\odot M_{T})^{\top}\right]^{\top},$ (3)
@@
where $\alpha_{t}\in\mathbb{R}$ are scalar weights (e.g., uniform or task-prioritized). The merged adapter becomes:
$\Delta_{\text{merge}}=A^{\prime}B^{\prime}=\sum_{t=1}^{T}\alpha_{t}A_{t}(B_{t}\odot M_{t})=\sum_{t=1}^{T}\alpha_{t}\Delta_{t}.$ (4)
This method ensures that the contribution of each adapter is preserved and interpretable, with minimal interference due to the approximate orthogonality of the LoRIs.
Linear Merging (Task Arithmetic). Alternatively, weighted sums of $A_{t}$ and $B_{t}\odot M_{t}$ are computed independently and then multiplied to produce the merged adapter:
$\Delta_{\text{merge}}=\left(\sum_{t=1}^{T}\alpha_{t}A_{t}\right)\left(\sum_{t=1}^{T}\alpha_{t}(B_{t}\odot M_{t})\right)=\sum_{s=1}^{T}\sum_{t=1}^{T}\alpha_{s}\alpha_{t}A_{s}(B_{t}\odot M_{t}).$ (5)
Linear merging introduces the cross terms $A_{s}(B_{t}\odot M_{t})$ for $s\neq t$, as each $A_{s}$ interacts with all $B_{t}\odot M_{t}$. These interactions can result in small but non-negligible interference. However, the overall interference remains bounded and is typically minor. In practice, concatenated merging is more effective at reducing cross-task interference than linear merging.
@@ 2.3 Reducing Interference in Continual Learning via Sparsity
In safety-critical applications, each newly introduced adapter must preserve the safety alignment of the base model, because we have to ensure that adaptations to new tasks do not compromise established safety behaviors. A straightforward solution is to merge the safety LoRI into the deployed model during every inference. However, as we will show in Section 3.4, this method may be insufficient for scenarios that demand strong safety guarantees. In such cases, as illustrated in Figure 2(c), a more reliable solution is to adopt a two-phase continual learning process for each LoRI adapter to reinforce safety:
@@
This method does not require recalibrating masks for each task or performing multiple rounds of continual learning. Notably, we do not enforce non-overlapping masks $M_{t}\cap M_{\text{safety}}=\emptyset$. Enforcing such a constraint would require recalibrating masks after the safety alignment phase due to the reduced parameter space, and could potentially degrade performance on downstream tasks. Empirically, we observe that mask overlap across tasks is minimal without explicitly enforcing non-overlap. At a 90% sparsity ratio, the average overlap between task-specific masks is $\sim 1\%$.
@@
Continual learning models are vulnerable to catastrophic forgetting (Li & Hoiem, 2017; Dong et al., 2023; Luo et al., 2023), where updates for new tasks can overwrite and degrade previously learned knowledge. Despite the minimal overlap between task-specific masks, the sparsity in $B_{t}$ induced by $M_{t}$ enables LoRI to facilitate isolated parameter updates for safety alignment and task adaptation. As a result, LoRI minimizes cross-task interference and mitigates catastrophic forgetting in safety alignment.

### v2

@@ 1 Introduction
Large language models (LLMs) (Brown et al., 2020; Touvron et al., 2023; Chowdhery et al., 2023) have transformed deep learning, showcasing remarkable capabilities across various domains. However, their deployment remains computationally demanding, particularly when fine-tuning is required to adapt to downstream tasks or align with human preferences. To mitigate the high resource costs, researchers have developed a range of parameter-efficient fine-tuning (PEFT) techniques. Among these techniques, LoRA (Hu et al., 2021) has gained widespread adoption due to its compelling balance of performance and efficiency. Nevertheless, LoRA still introduces notable memory overhead, particularly in large-scale models. Consequently, recent research has focused on further optimizing LoRA by reducing the number of trainable parameters without compromising performance (Kopiczko et al., 2023; Ding et al., 2023; Zhang et al., 2023b).
@@
Recent studies (Yu et al., 2024; Panda et al., 2024) have shown that delta parameters – the differences between fine-tuned and pretrained model weights – exhibit significant redundancy. Furthermore, previous works (Zhang et al., 2023b; Zhu et al., 2024) have observed that freezing matrices $A$ in LoRA often achieves comparable performance to training them. Motivated by these findings, we propose LoRA with Reduced Interference (LoRI). LoRI keeps matrices $A$ fixed as random projections, while training matrices $B$ using task-specific sparse masks. To retain the most critical elements of $B$, LoRI performs a calibration process to extract sparse masks by selecting the highest-magnitude elements across all layers and projections. As shown in Figure 1(a), LoRI maintains performance even with 90% sparsity in $B$ while keeping $A$ frozen. This demonstrates that adaptation does not require updating $A$, and that $B$ has considerable redundancy. By applying more constrained updates than LoRA, LoRI significantly reduces the number of trainable parameters while better preserving the pretrained model’s knowledge during adaptation.
@@
Figure 1: (a) Varying sparsity ratios in matrices $B$ while freezing $A$. Performance remains stable even at 90% sparsity in matrices $B$. (b) Merging three adapters via weighted averaging. LoRA suffers degradation due to parameter interference, while LoRI preserves task performance. (c) Continual learning from Safety to NLU. LoRA suffers from catastrophic forgetting, while LoRI retains safety alignment. Results for NLU are averaged over eight tasks. GSM8K accuracy (Math), HumanEval pass@10 (Code), and HEx-PHI refusal rate (Safety) are reported individually. Base model: Llama-3-8B, rank $r=32$.
@@
Multi-task learning is essential for enabling versatile models with multi-task capabilities, which is traditionally performed via joint training on a combination of task-specific datasets (Caruana, 1997; Sener & Koltun, 2018). However, training large models on this data mixture is prohibitively expensive in terms of time and compute. Model merging is a training-free alternative for building powerful models by combining existing ones (Ilharco et al., 2022; Yadav et al., 2023; Yu et al., 2024). This approach is well-suited for merging LoRA adapters, enabling multi-task capabilities within a single model during inference (Wang et al., 2024a; Prabhakar et al., 2024; Stoica et al., 2024). However, as shown in Figure 1(b), directly merging heterogeneous LoRAs often results in parameter interference, leading to degraded performance compared to single-task LoRAs. Additionally, many existing merging methods require trial-and-error to identify the optimal method for a specific combination of tasks. LoRI addresses these challenges by using fixed, randomly initialized projection $A$, which maps task-specific adapters into approximately orthogonal subspaces. This reduces interference when merging multiple adapters. In addition, LoRI enables adapter merging without manual selection of merging methods.
@@
Beyond multi-tasking, safety-critical scenarios require that each newly introduced adapter enhances model capabilities while preserving the safety alignment of the pretrained base model (Qi et al., 2023). LoRI provides a lightweight continual learning approach for adapting models while preserving safety, where training is performed sequentially across tasks (Lopez-Paz & Ranzato, 2017; Wu et al., 2022; Ouyang et al., 2022). The strategy involves first fine-tuning an adapter on safety data to establish alignment, followed by separate adaptation to each downstream task. However, as illustrated in Figure 1(c), continual learning often leads to catastrophic forgetting (Li & Hoiem, 2017; Dong et al., 2023; Luo et al., 2023), wherein the adaptation to new tasks substantially compromises previously acquired knowledge. LoRI mitigates forgetting by leveraging the sparsity of projection $B$ through task-specific masks. This isolation of parameter updates across tasks facilitates continual learning with minimal interference, preserving both safety and task effectiveness.
@@ 2.1 Freezing Low-Rank Projections with Sparse Masking
LoRA (Hu et al., 2021) fine-tunes a weight update matrix as a product of two low-rank matrices to adapt LLMs to new tasks. Formally, for a specific task $t$, given a pretrained weight matrix $W_{0}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$, the weight update $\Delta_{t}\in\mathbb{R}^{d_{\text{in}}\times d_{\text{out}}}$ is constrained to a low-rank decomposition:
$h=xW_{0}+x\Delta_{t}=xW_{0}+xA_{t}B_{t}.$ (1)
@@
Typically, the low-rank projection matrices $A_{t}$ and the low-rank expansion matrices $B_{t}$ are updated via gradient descent. Matrices $A_{t}$ are usually initialized with Kaiming Uniform distribution (He et al., 2015), while matrices $B_{t}$ are initialized to zero, ensuring that $\Delta_{t}=0$ at the start of training. However, in LoRI, we fix $A_{t}$ as random projections, meaning that the model only learns how to combine the fixed subspace via $B_{t}$. By freezing $A_{t}$, we eliminate the need to store their gradients and optimizer states, thereby reducing memory consumption. During inference, similar to LoRA, LoRI merges the low-rank updates by adding $A_{t}B_{t}$ to $W_{0}$, ensuring no additional inference latency compared to full fine-tuning.
@@
Figure 2: Overview of the proposed LoRI method. (a) LoRI freezes the projection matrices $A_{t}$ and sparsely updates $B_{t}$ using task-specific masks $M_{t}$. (b) LoRI enables adapter merging of multiple task-specific adapters with reduced parameter interference. (c) LoRI builds safety adapters by continual learning with reduced catastrophic forgetting.
@@
[…]
It is well established that the importance of projection matrices varies significantly across different layers and projections (Zhang et al., 2023a; d; Kopiczko et al., 2023). Our masking strategy enables global comparison of parameters and facilitates effective allocation of the parameter budget determined by the sparsity ratio. Notably, the masks for each task $t$ are calibrated only once and can be reused as needed.
@@ 2.2 Reducing Interference in Adapter Merging via Orthogonality
A central challenge in adapter merging is parameter interference, where combining multiple adapters leads to degraded performance due to conflicting parameter updates. Given a set of trained LoRI adapters $\{\Delta_{1},\Delta_{2},\dots,\Delta_{T}\}$, the goal is to construct a unified model that combines knowledge from all tasks with minimal interference, as illustrated in Figure 2(b). Formally, we define the excess loss due to parameter interference for a specific task $t$ as:
$\mathcal{I}_{t}=\mathcal{L}_{t}(W_{\text{merge}})-\mathcal{L}_{t}(W_{0}+\alpha_{t}\Delta_{t}),$ (3)
where $W_{\text{merge}}$ is the merged model, $W_{0}$ is the pretrained weight matrix, $\Delta_{t}$ is the LoRI adapter for task $t$, $\alpha_{t}\in\mathbb{R}$ is a scalar weight, and $\mathcal{L}_{t}$ is the loss function for task $t$. A high $\mathcal{I}_{t}$ indicates significant interference.
LoRI mitigates this interference by leveraging approximate orthogonality, achieved by freezing the projection matrices $A_{t}$ as independent random matrices. This design leads to the following property, whose proof is provided in Appendix C:
@@
Let $A_{s},A_{t}\in\mathbb{R}^{d_{\text{in}}\times r}$ be independent random matrices with i.i.d. entries drawn from a Kaiming Uniform distribution for distinct tasks $s\neq t$. Let their corresponding LoRI adapters be $\Delta_{s}=A_{s}(B_{s}\odot M_{s})$ and $\Delta_{t}=A_{t}(B_{t}\odot M_{t})$, where the trained matrices $(B_{s}\odot M_{s})$ and $(B_{t}\odot M_{t})$ have finite Frobenius norms. Under the condition that $r\ll d_{\text{in}}$, as the input dimension $d_{\text{in}}\to\infty$, the adapters are approximately orthogonal:
$\langle\Delta_{s},\Delta_{t}\rangle_{F}\to 0\quad\text{in probability}.$ (4)
@@
We describe two merging methods: concatenated merging (weighted averaging) and linear merging (Task Arithmetic) (Ilharco et al., 2022), both of which exploit the approximate orthogonality of LoRIs.
Concatenated Merging (Weighted Averaging).
This method constructs the merged model by creating a weighted sum of individual task adapters. This is achieved by concatenating the weighted $A$ and masked $B$ matrices:
$A^{\prime}=[\alpha_{1}A_{1}\ \alpha_{2}A_{2}\ \dots\ \alpha_{T}A_{T}],\quad B^{\prime}=\left[(B_{1}\odot M_{1})^{\top},\dots,(B_{T}\odot M_{T})^{\top}\right]^{\top},$ (5)
@@
where $\alpha_{t}\in\mathbb{R}$ are scalar weights (e.g., uniform or task-prioritized). The final merged model is then formed by adding their product to the base model weights:
$W_{\text{merge}}=W_{0}+A^{\prime}B^{\prime}=W_{0}+\sum_{t=1}^{T}\alpha_{t}A_{t}(B_{t}\odot M_{t})=W_{0}+\sum_{t=1}^{T}\alpha_{t}\Delta_{t}.$ (6)
By summing approximately orthogonal adapters, we ensure that the updates for each task occupy largely disjoint subspaces, thereby reducing interference (Ilharco et al., 2022; Ortiz-Jimenez et al., 2023; Xiong et al., 2024).
The reduction in interference can be explained by a theoretical sketch based on two key assumptions. The first is the local linearity of the loss landscape (Li et al., 2018), which allows for a first-order Taylor approximation. The second is the gradient alignment assumption, formally expressed as $\nabla\mathcal{L}_{t}(W_{0}+\alpha_{t}\Delta_{t})\propto\Delta_{t}$. This posits that at a task’s solution, the direction of steepest descent is primarily aligned with the adapter updates already made for that task. Under these assumptions, the excess loss $\mathcal{I}_{t}$ is approximately the inner product of the gradient and the updates from the other tasks:
$\mathcal{I}_{t}\approx\left\langle\nabla\mathcal{L}_{t}(W_{0}+\alpha_{t}\Delta_{t}),\sum_{s\neq t}\alpha_{s}\Delta_{s}\right\rangle_{F}\propto\sum_{s\neq t}\alpha_{k}\langle\Delta_{t},\Delta_{s}\rangle_{F}.$ (7)
Since Property 1 establishes that $\langle\Delta_{t},\Delta_{s}\rangle_{F}\to 0$ for $s\neq t$, the total interference loss becomes negligible: $\mathcal{I}_{t}\approx 0$. This heuristic argument provides strong intuition for why concatenated merging is effective, which is then validated by our empirical results.
Linear Merging (Task Arithmetic).
Alternatively, the merged model can be formed by summing the $A_{t}$ and masked $B_{t}$ matrices independently before multiplication:
$W_{\text{merge}}=W_{0}+\left(\sum_{t=1}^{T}\alpha_{t}A_{t}\right)\left(\sum_{t=1}^{T}\alpha_{t}(B_{t}\odot M_{t})\right)=W_{0}+\sum_{s=1}^{T}\sum_{t=1}^{T}\alpha_{s}\alpha_{t}A_{s}(B_{t}\odot M_{t}).$ (8)
While concatenated merging directly sums approximately orthogonal adapters, this linear merging approach introduces problematic cross-terms $\alpha_{s}\alpha_{t}A_{s}(B_{t}\odot M_{t})$ for $s\neq t$. These terms cause interference because components like $\{A_{s}(B_{t}\odot M_{t})\}_{t=1}^{T}$ for a fixed $s$ are generally not mutually orthogonal. As a result, concatenated merging offers a cleaner and empirically more effective strategy for combining LoRI adapters.
@@ 2.3 Reducing Interference in Continual Learning via Sparsity
For safety-critical applications, ensuring that new task adaptations do not compromise established safety behaviors is crucial. Therefore, each newly introduced adapter must preserve the base model’s safety alignment. A straightforward approach to achieve this is to merge a safety LoRI adapter into the deployed model during every inference. However, as we will show in Section 3.4, this method may be insufficient for scenarios that demand strong safety guarantees. In such cases, as illustrated in Figure 2(c), a more reliable solution is to adopt a two-phase continual learning process for each LoRI adapter to reinforce safety:
@@
[…]
@@
Continual learning models are vulnerable to catastrophic forgetting (Li & Hoiem, 2017; Dong et al., 2023; Luo et al., 2023), where updates for new tasks can overwrite and degrade previously learned knowledge. Despite the slight overlap between task-specific masks, the sparsity in $B_{t}$ induced by $M_{t}$ enables LoRI to facilitate isolated parameter updates for safety alignment and task adaptation. As a result, LoRI minimizes cross-task interference and mitigates catastrophic forgetting in safety alignment.